OpenStudy (anonymous):

Proving using induction Sum {i=0 to n} (n choose i) = 2^n for all n >= 0. Ok, so the idea is to add the next term. But what is the next term?

OpenStudy (anonymous):

k+1 ∑ (k + 1 choose i) ???? i=0

OpenStudy (kinggeorge):

That's exactly right. You just have to remember that one of the definitions for the binomial coefficients is\[\binom{n}{k}=\binom{n-1}{k}+\binom{n-1}{k-1}\]

OpenStudy (kc_kennylau):

Derived from the Pascal's triangle

OpenStudy (anonymous):

@kropot72 O.o is that really the definition? or a derived formula?

OpenStudy (kc_kennylau):

Well how do you calculate one grid in the Pascal Triangle?

OpenStudy (anonymous):

@kc_kennylau add the number from the left to the number from the right?

OpenStudy (kc_kennylau):

That's exactly what the formula is doing

OpenStudy (kinggeorge):

If you say that \[\binom{n}{0}=\binom{n}{n}=1\]for all positive integers \(n\), it's a perfectly valid recursive definition.

OpenStudy (kinggeorge):

And the definition can be thought to be exactly equivalent to the construction of Pascal's triangle.

OpenStudy (anonymous):

ahh I see. I was used to thinking (nCk) = n! / ( (n-k)! k! )

OpenStudy (anonymous):

so k+1 ∑ (k + 1 choose i) is correct? i=0

OpenStudy (anonymous):

wait no i don't think it is O.O

OpenStudy (kinggeorge):

You want to show that that sum is equal to \(2^{k+1}\) assuming \(\sum_{i=0}^k\binom{k}{i}=2^k\). If you've already shown it for \(n=0\), that would be enough finish the problem by induction.

OpenStudy (anonymous):

but n = 0 is the base case. What about the inductive step? That's what i'm stucked on. I know the assumption is that (kC0) + (kC1) + ... (kCk) = 2^k for some integer k, now i need to add the next term, which is (k+1 C k) ??

OpenStudy (anonymous):

or is it (k+1 C i) ?

OpenStudy (kinggeorge):

The inductive step is assuming that\[\sum_{i=0}^k\binom{k}{i}=2^k.\]You want to show that\[\sum_{i=0}^{k+1}\binom{k+1}{i}=2^{k+1}\]somehow using the inductive hypothesis. The trick to do this, is to write each term as a sum of smaller terms using the definition of the binomial coefficients I gave above. So the first step would be\[\sum_{i=0}^{k+1}\binom{k+1}{i}=\sum_{i=0}^{k+1}\binom{k}{i}+\binom{k}{i-1}.\]You need to take a bit of care with the indices, but from this point, you should be able to apply the inductive hypothesis without too much trouble.

OpenStudy (anonymous):

ahh I see. The "next term" in this context is the sum of the coefficient of next row in the pascal triangle. I just thought it's the next term on the current k row -.- No wonder i couldn't reach 2^(k+1). Thank you ^.^ @KingGeorge

OpenStudy (kinggeorge):

This is a nice example of what could be considered a non-standard inductive proof for exactly the reason. There's no real "next term" to add on, so you have to be a little creative with what happens when you increment \(k\).